OpenStudy (anonymous):

Use a graph to find the range of the function f(x) = x^2 - 9 on the domain -3 < x < 4

OpenStudy (anonymous):

Domain is :\[-3 \le x \le 4\]

OpenStudy (tkhunny):

Well,I guess you should graph it and look?

OpenStudy (anonymous):

but i dont know how to get the answer

OpenStudy (freckles):

Are you saying you don't know how to graph y=x^2-9? Try to compare the this: The following is a parabola y=a(x-h)^2+k where (h,k) is vertex of the parabola

OpenStudy (anonymous):

ok thanks